Suppose that your company is expected to pay a dividend of $1.50 per share next year. There has been a steady growth in dividends of 5.1% per year and the market expects that to continue. What is the current price of the stock if the required return is 10%?

Price=
Answer #1
Price = $30.61
Statement showing Computations
Particulars Amount
P0 = D1/(ke-g)
Where P0 is the price
Ke is the required return 10.00%
g is the growth rate 5.10%
D1 is dividend at end of year                     1.50
P0 = 1.5/(10% - 5.1%)
P0 = 1.5/4.9%
P0 = $30.61